OpenStudy (anonymous):

The formula for the volume of a sphere is V= 4/3 (PI)r^3 . Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3

OpenStudy (anonymous):

15in^3=(pi)r^3 15*(3/4)/pi=r^3 3.581in^3=r^3 r-1.5299in r-1.53in

OpenStudy (anonymous):

@uri